OpenStudy (anonymous):

Point A is located at (0,4) and point B is located at (-2,-3). Find the point that is 1/4 the distance from point A to point B A. (-1, 0.75) B. (-0.75, 2) C. (-0.5, 2.25) D. (-0.25, 3)

OpenStudy (anonymous):

@nirmalnema

OpenStudy (lxelle):

Find the midpoint, M It's the average of x & y for A & B. M is (-1,1/2) Then find Q, midway between A & M.

OpenStudy (nirmalnema):

use section formula.. let that point be(x,y) let m/n=1/4 x=(mx2+nx1)/m+n y=(my2+ny1)/m+n

OpenStudy (anonymous):

@nirmalnema I dont understand

OpenStudy (nirmalnema):

which part ?? u dont understand

OpenStudy (anonymous):

@nirmalnema the entire thing lol

OpenStudy (nirmalnema):

you need to find that point which divides the distance b/w given points in the ratio1/4

OpenStudy (nirmalnema):

now do u understand.??